OpenStudy (anonymous):

Please help me with these steps!! Will give medal! Complete the table below to solve the equation 2.5x – 10.5 = 64(0.5x). x f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 2 3 4 5 6

OpenStudy (anonymous):

This is a table graph, and I have to solve both equations separately!

OpenStudy (anonymous):

@ganeshie8 @mathstudent55 @Abhisar @thomaster @.Sam. @dan815 @Compassionate @uri @Preetha @sammixboo @Loser66 Someone please help!

jhonyy9 (jhonyy9):

so when there are for example f(x)=2x+1 and you need calculi f(1) what you need to do ?

jhonyy9 (jhonyy9):

f(1) mean that x=1 yes ?

OpenStudy (anonymous):

I dont understand math that well...

OpenStudy (anonymous):

but yes

jhonyy9 (jhonyy9):

so than f(1)=2*1+1 so f(1)= 2+1 so f(1)=3 ok ?

OpenStudy (anonymous):

Ok

jhonyy9 (jhonyy9):

so just you need using this example in case of your exercise

jhonyy9 (jhonyy9):

1. f(x)=2,5x -10,5 x=2 so this mean f(2)=2,5*2-10 so f(2)=5-10 so f(2)=-5 rigth ?

OpenStudy (anonymous):

X f(x) = 2.5x – 10.5 g(x) = 64(0.5x) 1.F(x) = 2.5x – 10.5 1.g(x) = 64(0.5) 2.2.5x – 10.5 2.64(0.5x) 3.2.5x – 10.5 3.32x 4.2.5x – 32x 4.10.5 5.–29.5x 5.10.5 6.x = -10.5 6.29.5

OpenStudy (anonymous):

Would it look like this??

jhonyy9 (jhonyy9):

sorry but from these i dont understand nothing

OpenStudy (anonymous):

Sorry, something is going on with the messenger thing! Hold on please

jhonyy9 (jhonyy9):

x f(x)=2,5x-10 g(x)=64(0,5x) so g(x)=32x 2 f(2)=2,5*2-10 = 5 -10 = - 5 g(2)=32*2=64 3 f(3)=2,5*3-10=7,5-10 = -2,5 g(3)=32*3=96 4 f(4)= g(4)= 5 f(5)= g(5)= 6 f(6)= g(6)= after these hope so much that you can calculi them easy ,yes ?

jhonyy9 (jhonyy9):

ok ?

jhonyy9 (jhonyy9):

f(4)= ???

jhonyy9 (jhonyy9):

do you understand it now ?

OpenStudy (anonymous):

Ok, for the second on (g(x)) would it continue like this 128, 160, and 192? and for the first it would be -5.5, -3, -0.5, 2, 4.5?

jhonyy9 (jhonyy9):

for the second are rigth but for the first how you have got these results ?

OpenStudy (anonymous):

Is that right?

jhonyy9 (jhonyy9):

the first and second result for the first function i have wrote above so you need continue with f(4)= ?

OpenStudy (anonymous):

But with what I wrote is that right or no?

jhonyy9 (jhonyy9):

ohh sorry i have missed 0,5 because there is 10,5 not just 10 so than f(2)=2,5*2-10,5 = 5-10,5 = -5.5 f(3)=7.5-10,5=-3 f(4)=10-10,5=-0,5 f(5)=12,5-10,5=2 f(6)=13,5-10,5=3 sorry these will be right sure

OpenStudy (anonymous):

Ok great! Thanks a bunch! And the second one is right correct?

jhonyy9 (jhonyy9):

yes is right sure good luck was my pleasure bye